The length of a rectangle is 8cm longer than the width. If the width is increased by 4cm and the length is decreased by 5cm, the area of the rectangle remains the same. Find the original dimensions of the rectangle.
Answer by ikleyn(52781)

.
Let w be the width.

Then your equation is

w*(w+8) = (w+4)*((w+8)-5),    or, which is the same,

w*(w+8) = (w+4)*(w+3),

w^2 + 8w = w^2 + 4w + 3w + 12,

w = 12.


Check.  12*(12+8) = 240.

        (12+4)(12+3) = 16*15 = 240.  ! Correct !


Answer.  the original dimensions of the rectangle are 12 x 8 cm.

Solved.